OpenStudy (anonymous):

Order -3, 2, and -5 from least to greatest. a. -5, -3, 2 c. 2, -3, -5 b. -5, 2, -3 d. -3, -2, -5

OpenStudy (anonymous):

its A

OpenStudy (anonymous):

a

OpenStudy (anonymous):

The answer is A.

OpenStudy (anonymous):

Simplify 10x – 5 + 4x + 2(x – 7). a. 16x – 2 c. 16x – 10 b. 16x – 12 d. 16x – 19

OpenStudy (dmezzullo):

16x??????

OpenStudy (dmezzullo):

should be 14 or 15x

OpenStudy (anonymous):

D

OpenStudy (dmezzullo):

its b

OpenStudy (anonymous):

16xFind the slope of the line that passes through (-10, 5) and (2, 5). a. -8 c. 0 b. undefined d. -12

OpenStudy (dmezzullo):

cse -5+2=-3+7=10